Calculate the thrust of a steel ball of 10cm radius immersed in water. (Archimedes Principle) Show solution.

1
Expert's answer
2021-12-07T20:11:46-0500

The volume of the ball:


"V=\frac 43\pi r^3=4.19·10^{-3}\text{ m}^3."

The buoyancy force:


"F=\rho gV=41.1\text{ N}."